OpenStudy (anonymous):

Select the statement that best justifies the conclusion based on the given information. If a(b+c)=d, then ab+ac=d Associative Commutative Distributive Closure

OpenStudy (sleepyjess):

Hey there! Welcome to OpenStudy. I will give you an example of each of the answers and you can decide with the information I give you. Associative-4+(7+2)=(4+7)+2 Commutative-5*3*9=5*9*3=9*5*3 Distributive-2(x+y)=2x+2y Closure-Says that 2 real numbers added together equals a real number. 4+5=9

OpenStudy (anonymous):

Thanks. That will help with my other questions as well.

OpenStudy (sleepyjess):

No problem!